# Thermal Inertia Controls on Titan’s Surface Temperature and Planetary Boundary Layer Structure

### v1

@@ Abstract
Understanding Titan’s planetary boundary layer (PBL)—the lowest region of the atmosphere influenced by surface conditions—remains challenging due to Titan’s thick atmosphere and limited observations. Previous modeling studies have also produced inconsistent estimates of surface temperature, a critical determinant of PBL behavior, often without clear explanations grounded in surface energy balance. In this study, we develop a theoretical framework and apply a three-dimensional dry general circulation model (GCM) to investigate how surface thermal inertia influences surface energy balance and temperature variability across diurnal and seasonal timescales. At diurnal timescales, lower thermal inertia surfaces exhibit larger temperature swings and enhanced sensible heat fluxes due to inefficient subsurface heat conduction. In contrast, at seasonal timescales, surface temperature variations show weak sensitivity to thermal inertia, as atmospheric damping tends to dominate over subsurface conduction. The PBL depth ranges from a few hundred meters to 1,000 [m] on diurnal timescales, while seasonal maxima reach 2,000–3,000 [m], supporting the interpretation from a previous study that the Huygens probe captured the two PBL structures. Simulated seasonal winds at the Huygens landing site successfully reproduce key observed features, including near-surface retrograde winds and meridional wind reversals within the lowest few kilometers, consistent with Titan’s cross-equatorial Hadley circulation. Simulations at the planned Dragonfly landing site predict shallower thermal PBLs with broadly similar wind patterns. This work establishes a physically grounded framework for understanding Titan’s surface temperature and boundary layer variability, and offers a unified explanation of Titan’s PBL behavior that provides improved guidance for future missions.
@@ 1 Introduction
The planetary boundary layer (PBL) is the lowest atmospheric layer on terrestrial worlds, and is characterized by significant fluxes of heat, moisture, and momentum between the surface and atmosphere, driven by turbulent motions. On Earth, the PBL undergoes substantial diurnal and seasonal variations due to surface heating and convection. These transitions between temperature extrema, whether diurnal or seasonal, generate distinct atmospheric stability regimes that influence critical weather and climate processes, including cloud formation, precipitation, local breezes, and thunderstorms (Stull, 1988; Garratt, 1992; Lee, 2016). The phenomenon of the PBL extends beyond Earth to other celestial bodies possessing both a solid surface and an atmosphere (Petrosyan et al., 2011; Lebonnois et al., 2013; Flasar et al., 2014). As such, studying the PBL of other planetary bodies provides valuable insights into general PBL processes, making it crucial for comparative planetology. Mars, for instance, exhibits a PBL similar to Earth’s, where convection plays a fundamental role (Hinson et al., 2010; Spiga et al., 2010).
@@
Titan, Saturn’s largest satellite, presents unique geological landscapes with dry, vast linear dunes at the equator contrasting with lakes and seas in the polar regions (Lorenz et al., 2006; Aharonson et al., 2014; Hayes, 2016). These striking geographical contrasts make understanding Titan’s PBL essential for explaining the formation and evolution of such diverse landscapes. Due to Titan’s vast distance from the Sun and its thick, opaque atmosphere, average surface insolation is estimated at approximately 0.1 % of Earth’s, around 1.0 [Wm-2] (McKay et al., 1989, 1991; Tomasko et al., 2008). However, given Titan’s extended day (16 Earth days) and year (29.5 Earth years), the accumulated surface energy over these cycles may suffice to generate notable surface temperature variations and PBL responses. Observational evidence remains limited, especially for a diurnal cycle, with only one in-situ vertical thermal profile near the equatorial surface from the Huygens landing (Fulchignoni et al., 2005) and scattered radio occultation profiles across the moon (Flasar et al., 1981; Lindal et al., 1983; Samuelson et al., 1997; Courtin & Kim, 2002; Schinder et al., 2011, 2012, 2020). Consequently, modeling studies are vital to supplementing the limited observational data and revealing whether Titan’s PBL responds to surface heating through convective processes in ways analogous to Earth and Mars, both diurnally and seasonally.
@@
Understanding how surface temperature responds to solar forcing is essential for interpreting variations in the PBL. Within the context of modeling surface temperature variations on terrestrial planets, thermal inertia, defined as [JmsK-1] (hereafter we will use thermal inertia units; TIU), serves as a crucial parameter for quantifying a surface’s thermal response to energy input (Price, 1977; Putzig et al., 2005; MacKenzie et al., 2019). Here, represents the density [kgm-3], denotes the specific heat capacity [JkgK-1], and is the thermal conductivity [WmK-1] of the surface. Materials with lower thermal inertia tend to exhibit larger temperature variations, reflecting their diminished capacity to store and conduct heat (Stull, 1988; Garratt, 1992).
@@
Using a dry general circulation model (GCM) of Titan’s atmosphere coupled with a simplified subsurface scheme, Tokano (2005) reported diurnal and seasonal surface temperature variations of a few kelvins near the equator—close to the latitude of the Huygens landing at 10.3°S—assuming a uniform thermal inertia of [TIU], characteristic of porous ice regolith. Williams et al. (2012) employed a one-dimensional energy-balance model, considering the Huygens landing site with a fixed near-surface atmospheric temperature, to investigate surface temperature variations. Their results indicated smaller diurnal and seasonal surface temperature variations, with amplitudes less than 0.1 [K] under dry conditions and a thermal inertia of [TIU], contradicting the findings of Tokano (2005). Subsequently, Charnay & Lebonnois (2012) reported diurnal fluctuations around 0.3 [K] in surface temperatures at the Huygens landing site using a dry GCM with [TIU] (Lebonnois et al., 2012). They proposed that the Huygens measurement of a near-adiabatic layer reaching 300 [m] (Tokano et al., 2006) reflected ongoing convection during the probe’s descent (at approximately 10 am local time), while changes in the lapse rate at 800 [m] and 2,000 [m] corresponded to residual signatures of maximum diurnal PBL extension and seasonal Hadley circulation effects, respectively. More recently, MacKenzie et al. (2019) implemented a global thermal inertia map in a dry GCM (Lora et al., 2015), demonstrating that seasonal surface temperature patterns show limited sensitivity to local thermal inertia values.
@@
Building on these insights, this study develops a theoretical framework that links surface temperature variations to thermal inertia and surface energy balance, and conducts dry GCM simulations to investigate how varying thermal inertia influences surface temperature and PBL behavior on Titan across diurnal and seasonal timescales. Section 2 introduces the theoretical foundations governing surface temperature variability and outlines the modeling approach. Section 3 presents GCM simulation results for surface temperature variations and demonstrates how the theoretical framework explains them at both diurnal and seasonal timescales. Section 4 examines PBL structures from GCM simulations and their relationship to surface temperatures. Section 5 evaluates expected conditions at the Dragonfly landing site. Section 6 discusses the implications of the findings and and concludes by highlighting the distinct roles of atmospheric damping and subsurface heat conduction in regulating surface temperature and PBL structures—dominant at seasonal and diurnal timescales, respectively. Overall, this study provides a critical reassessment of the atmospheric profile measured by Huygens and offers new insights into Titan’s lower atmosphere, informing expectations for the upcoming Dragonfly mission.
@@ 2.1 Theory
where and [Wm-2] represent the net radiative flux into and the sensible heat flux out of the surface, respectively. The residual, [Wm-2], is the net energy flux at the surface available for subsurface heat conduction.
@@
where [JmK-1] collectively denotes volumetric heat capacity. Here, homogeneous subsurface properties are considered for simplicity and to compare with previous results. The net energy flux at the surface can be simplified as a single harmonic oscillator with linear atmospheric damping (e.g., Dickinson, 1988):
where [Wm-2] is the amplitude of the periodic flux, and [Wm-2K-1] is the atmospheric damping coefficient. Over land surfaces, as assumed in dry GCM configurations, an additional lower boundary condition must be applied to the subsurface; for example:
where [K] is some mean temperature in the deep subsurface, and [m] is the thermal skin depth corresponding to a forcing frequency of [s-1], whether it be diurnal or annual. The magnitude of surface temperature fluctuations is expressed as (see Appendix A for the full derivation):
where [Wm-2] denotes the amplitude of fluctuations in . The thermal inertia [TIU], combined with [s-0.5] represents damping of the energy ([Wm-2K-1]) received at the surface, due to subsurface heat conduction. Meanwhile, [Wm-2K-1], which is the magnitude of , accounts for damping due to direct energy exchange between the surface and the atmosphere.
For Titan, the temperature of the overlying atmosphere, [K], can be reasonably assumed to remain constant due to its long radiative timescale (Flasar et al., 1981, 2014). Notably, the linear atmospheric damping term in Eq. (4) takes the same form as the expression for sensible heat flux:
where [kgm-3] is the density of Titan’s lower atmosphere as measured by the Huygens probe, is the specific heat capacity of the N2-dominated atmosphere [JkgK-1], [dimensionless] is the surface heat transfer coefficient used in GCMs, and [m s-1] is the typical magnitude of the near-surface horizontal winds in GCM simulations (e.g., Lora et al., 2019). In the following sections, we examine how the surface temperature variations reproduced in our GCM simulations at diurnal and seasonal timescales correspond to the theoretical predictions, Eq (6).
@@ 2.2 Titan Atmospheric Model
Our investigation encompasses four simulation cases (Table LABEL:tab:table1). Cases 1, 2, and 3 employ a uniform volumetric heat capacity of [JmK-1], with varying thermal conductivities of , , and [WmK-1], respectively, corresponding to thermal inertia values of , , and [TIU]. Case 4 adopts a spatially varying thermal inertia map based on the “convective lakes” scenario of MacKenzie et al. (2019), which characterizes Titan’s surface using Cassini RADAR observations and classifies it into five terrain types: dunes, hummocky, labyrinth, plains, and lakes, which are assumed to be convecting. In our study, we fix across Cases 1–3 as the volumetric heat capacity values across terrain types in MacKenzie et al. (2019) are relatively similar in magnitude. In contrast, thermal conductivity spans over two orders of magnitude, and our three selected values are chosen to broadly represent this range.
The simulations are run at T21 horizontal resolution (64 × 32 longitude-latitude grid) with 50 atmospheric vertical levels and 15 subsurface vertical levels, using a 600 [s] time step. A shortwave albedo of 0.2 and surface thermal emissivity of 0.95 are adopted. The atmospheric vertical resolution in the lower few kilometers is maintained at several tens of meters to sufficiently resolve boundary layer variations. Subsurface layer depths follow and , where represents the diurnal skin depth for Titan and denotes the subsurface level. This configuration ensures adequate resolution for both diurnal and seasonal subsurface heat conduction by progressively increasing layer depth, with 5 layers within , 10 layers within (the annual skin depth), and 5 layers extending to the deepest region at .
The model is initialized from a cold start with uniform surface and subsurface temperatures of [K], and integrated for 200 Titan years with diurnally averaged output. For diurnal analyses, the simulation is extended by one additional Titan year with output saved every 16 Earth hours (i.e., 1/24 of a Titan day), focusing on a 30 Titan-day period centered around southern midsummer in the final year.
@@ 3.1 Diurnal Variations
Figure 1 presents the diurnal evolution of surface energy balance components and surface temperature near the Huygens landing site during midsummer for the four simulation cases. In Case 4, the plains terrain type ( [TIU]) is assumed at the Huygens landing site (MacKenzie et al., 2019). In Case 1, the net shortwave (SW) flux increases from 0.0 to approximately 1.3 [Wm-2] from nighttime to local noon, while the sensible heat flux peaks near 1.2 [Wm-2] around midday, indicating that a substantial portion of incoming energy is returned to the atmosphere. This behavior reflects the limited capacity of low thermal inertia surfaces to conduct heat into the subsurface, necessitating that excess surface energy be dissipated through sensible heat flux. As thermal inertia increases (Cases 2 and 3), the amplitude of the sensible heat flux correspondingly decreases (1.0 and 0.7 [Wm-2]), as subsurface heat conduction becomes more effective at removing energy from the surface. Meanwhile, the net infrared (IR) radiative flux remains relatively constant at approximately 0.07 [Wm-2] across all cases, consistent with the relatively small changes in surface temperature.
@@
The simulated diurnal surface temperature variations are approximately , , and [K] for Cases 1, 2, and 3, respectively. These values represent the mean and one standard deviation computed over a 30 Titan-day period centered on midsummer in the final simulation year. For a diurnal forcing frequency of [s-1] and net energy flux amplitudes of , , and [Wm-2] for Cases 1, 2, and 3, respectively—derived using the same 30 Titan-day averaging method as for the simulated surface temperature amplitudes—and with an atmospheric damping coefficient of [WmK-1], obtained as the 30-day mean value, the theoretical temperature variations according to Eqs. (6) and (7) are estimated to be , 0.16, and 0.13 [K]. These estimates closely align with the simulated values, reinforcing the validity of the theoretical framework despite its simplification of the net energy flux as a sine-like function. For Case 4, the [Wm-2] yields a theoretical expectation of [K], closely matching the [K] as well.
@@
Despite the fact that the net energy flux—the residual of the surface energy budget—increases with thermal inertia, the amplitude of surface temperature variations decreases. The decreasing amplitude of temperature variation arises because increasing thermal inertia enhances subsurface heat conduction, effectively buffering the surface temperature despite the higher available energy. The values of corresponding to the thermal inertias used in Cases 1, 2, and 3 are approximately 0.37, 1.17, and 3.69 [WmK-1], respectively. These values significantly exceed the atmospheric damping coefficient, [WmK-1], indicating that subsurface heat conduction is the dominant process damping surface temperature variations at diurnal timescales.

### v2

@@ Abstract
Understanding Titan’s planetary boundary layer (PBL)—the lowest region of the atmosphere influenced by surface conditions—remains challenging due to Titan’s thick atmosphere and limited observations. Previous modeling studies have also produced inconsistent estimates of surface temperature variability, a critical determinant of PBL behavior, often without clear explanations grounded in surface energy balance. Here, we develop a theoretical framework and apply a three-dimensional dry general circulation model (GCM) to investigate how surface thermal inertia influences surface energy balance and temperature variability across diurnal and seasonal timescales. At diurnal timescales, lower thermal inertia surfaces experience larger temperature fluctuations and enhanced daytime sensible heat fluxes due to less efficient subsurface heat conduction. In contrast, at seasonal timescales, surface temperature variations show weak sensitivity to thermal inertia, as atmospheric damping tends to dominate over subsurface conduction. The PBL depth ranges from a few hundred meters to 1,000 m on diurnal timescales, while seasonal maxima reach 2,000 m, supporting the interpretation from a previous study that the Huygens probe captured the two PBL structures. Simulated seasonal winds at the Huygens landing site successfully reproduce key observed features, including near-surface retrograde winds and meridional wind reversals within the lowest few kilometers, consistent with Titan’s cross-equatorial Hadley circulation. Simulations for the planned Dragonfly landing site predict shallower thermal PBLs with smaller fluctuation amplitudes, while maintaining similar wind patterns. This work establishes a physically grounded framework for understanding Titan’s surface temperature and boundary layer variability, and offers a unified explanation of Titan’s PBL behavior that provides improved guidance for future missions.
@@ 1 Introduction
The planetary boundary layer (PBL) is the lowest atmospheric layer on terrestrial worlds, and is characterized by significant fluxes of heat, moisture, and momentum between the surface and atmosphere, driven by turbulent motions. On Earth, the PBL undergoes substantial diurnal and seasonal variations due to surface heating and convection. These transitions between temperature extrema, whether diurnal or seasonal, generate distinct atmospheric stability regimes that influence critical weather and climate processes, including cloud formation, precipitation, local breezes, and thunderstorms (Stull, 1988; Garratt, 1992; Lee, 2016). The phenomenon of the PBL extends beyond Earth to other celestial bodies possessing both a solid surface and an atmosphere (Petrosyan et al., 2011; Flasar et al., 2014; Lebonnois et al., 2018). As such, studying the PBL of other planetary bodies provides valuable insights into general PBL processes, making it crucial for comparative planetology. Mars, for instance, exhibits a PBL similar to Earth’s, where convection plays a fundamental role (Hinson et al., 2010; Spiga et al., 2010).
@@
Titan, Saturn’s largest satellite, presents unique geological landscapes with dry, vast linear dunes at the equator contrasting with lakes and seas in the polar regions (Lorenz et al., 2006; Aharonson et al., 2014; Hayes, 2016). These striking geographical contrasts make understanding Titan’s PBL essential for explaining the formation and evolution of such diverse landscapes. Due to Titan’s vast distance from the Sun and its thick, opaque atmosphere, average surface insolation is estimated at approximately 0.1 % of Earth’s, around 1.0 Wm-2 (McKay et al., 1989, 1991; Tomasko et al., 2008). However, given Titan’s extended day (16 Earth days) and year (29.5 Earth years), the accumulated surface energy over these cycles may suffice to generate notable surface temperature variations and PBL responses. Observational evidence remains limited, especially for a diurnal cycle, with only one in-situ vertical thermal profile near the equatorial surface from the Huygens landing (Fulchignoni et al., 2005) and scattered radio occultation profiles across the moon (Flasar et al., 1981; Lindal et al., 1983; Samuelson et al., 1997; Courtin & Kim, 2002; Schinder et al., 2011, 2012, 2020). Consequently, modeling studies are vital to supplementing the limited observational data and revealing whether Titan’s PBL responds to surface heating through convective processes in ways analogous to Earth and Mars, both diurnally and seasonally.
@@
Understanding how surface temperature responds to solar forcing is essential for interpreting variations in the PBL. Within the context of modeling surface temperature variations on terrestrial planets, thermal inertia, defined as (JmsK-1; hereafter we will use thermal inertia units; TIU), serves as a crucial parameter for quantifying a surface’s thermal response to energy input (Price, 1977; Putzig et al., 2005; MacKenzie et al., 2019). Here, represents the density (kgm-3), denotes the specific heat capacity (JkgK-1), and is the thermal conductivity (WmK-1) of the surface. Materials with lower thermal inertia tend to exhibit larger temperature variations, reflecting their diminished capacity to store and conduct heat (Stull, 1988; Garratt, 1992).
@@
Using a dry general circulation model (GCM) of Titan’s atmosphere coupled with a simplified subsurface scheme, Tokano (2005) reported diurnal and seasonal surface temperature variations of a few kelvins near the equator—close to the latitude of the Huygens landing at 10.3°S—assuming a uniform thermal inertia of TIU, characteristic of porous ice regolith. Williams et al. (2012) employed a one-dimensional energy-balance model, considering the Huygens landing site with a fixed near-surface atmospheric temperature, to investigate surface temperature variations. Their results indicated smaller diurnal and seasonal surface temperature variations, with amplitudes less than 0.1 K under dry conditions and a thermal inertia of TIU, contradicting the findings of Tokano (2005). Subsequently, Charnay & Lebonnois (2012) reported diurnal fluctuations around 0.3 K in surface temperatures at the Huygens landing site using a dry GCM with TIU (Lebonnois et al., 2012). They proposed that the Huygens measurement of a near-adiabatic layer reaching 300 m (Tokano et al., 2006) reflected ongoing convection during the probe’s descent (at approximately 10 am local time), while changes in the lapse rate at 800 m and 2,000 m corresponded to residual signatures of maximum diurnal PBL extension and seasonal Hadley circulation effects, respectively. More recently, MacKenzie et al. (2019) implemented a global thermal inertia map in a dry GCM (Lora et al., 2015), demonstrating that seasonal surface temperature patterns show limited sensitivity to local thermal inertia values.
@@
Building on these insights, this study develops a theoretical framework that links surface temperature variations to thermal inertia and surface energy balance, and conducts dry GCM simulations to investigate how varying thermal inertia influences surface temperature and PBL behavior on Titan across diurnal and seasonal timescales. Section 2 introduces the theoretical foundations governing surface temperature variability and outlines the modeling approach. Section 3 presents GCM simulation results for surface temperature variations and demonstrates how the theoretical framework explains them at both diurnal and seasonal timescales. Section 4 examines PBL structures from GCM simulations and their relationship to surface temperatures. Section 5 evaluates expected conditions at the Dragonfly landing site. Section 6 discusses the implications of the findings and concludes by highlighting the distinct roles of atmospheric damping and subsurface heat conduction in regulating surface temperature and PBL structures—dominant at seasonal and diurnal timescales, respectively. Overall, this study provides a critical reassessment of the atmospheric profile measured by Huygens and offers new insights into Titan’s lower atmosphere, informing expectations for the upcoming Dragonfly mission.
@@ 2.1 Theory
where and (Wm-2) represent the net radiative flux into and the sensible heat flux out of the surface, respectively. The residual, (Wm-2), is the net energy flux at the surface available for subsurface heat conduction.
@@
where (JmK-1) collectively denotes volumetric heat capacity. Here, homogeneous subsurface properties are considered for simplicity and to compare with previous results. The net energy flux at the surface can be simplified as a single harmonic oscillator with linear atmospheric damping (e.g., Dickinson, 1988):
where is the amplitude of the periodic flux, and (Wm-2K-1) is the atmospheric damping coefficient. Over land surfaces, as assumed in dry GCM configurations, an additional lower boundary condition must be applied to the subsurface; for example:
where is some mean temperature in the deep subsurface, and (m) is the thermal skin depth corresponding to a forcing frequency of (s-1), whether it be diurnal or annual. The magnitude of surface temperature fluctuations is expressed as (see Appendix A for the full derivation):
where denotes the amplitude of fluctuations in . The thermal inertia , combined with represents damping of the energy (Wm-2K-1) received at the surface, due to subsurface heat conduction. Meanwhile, , which is the magnitude of , accounts for damping due to direct energy exchange between the surface and the atmosphere.
For Titan, the temperature of the overlying atmosphere, , can be reasonably assumed to remain constant due to its long radiative timescale (Flasar et al., 1981, 2014). Notably, the linear atmospheric damping term in Eq. (4) takes the same form as the expression for sensible heat flux:
where (kgm-3) is the density of Titan’s lower atmosphere as measured by the Huygens probe, is the specific heat capacity of the N2-dominated atmosphere (JkgK-1), is the surface heat transfer coefficient used in GCMs, and (m s-1) is the typical magnitude of the near-surface horizontal winds in GCM simulations (e.g., Lora et al., 2019). In the following sections, we examine how the surface temperature variations reproduced in our GCM simulations at diurnal and seasonal timescales correspond to the theoretical predictions, Eq (6).
@@ 2.2 Titan Atmospheric Model
Our investigation encompasses four simulation cases (Table LABEL:tab:table1). Cases 1, 2, and 3 employ a uniform volumetric heat capacity of JmK-1, with varying thermal conductivities of , , and WmK-1, respectively, corresponding to thermal inertia values of , , and TIU. Case 4 adopts a spatially varying thermal inertia map based on the “convective lakes” scenario of MacKenzie et al. (2019), which characterizes Titan’s surface using Cassini RADAR observations and classifies it into five terrain types: dunes, hummocky, labyrinth, plains, and lakes, which are assumed to be convecting. In our study, we fix across Cases 1–3 as the volumetric heat capacity values across terrain types in MacKenzie et al. (2019) are relatively similar in magnitude. In contrast, thermal conductivity spans over two orders of magnitude, and our three selected values are chosen to broadly represent this range.
The simulations are run at T21 horizontal resolution (64 × 32 longitude-latitude grid) with 50 atmospheric vertical levels and 15 subsurface vertical levels, using a 600 s time step. A shortwave albedo of 0.2 and surface thermal emissivity of 0.95 are adopted. The radiative transfer parameterization in TAM, which uses plane-parallel two stream approximations, is non-gray and accounts for multiple scattering, employing absorption coefficients and haze optical properties from observations (Lora et al., 2015) and producing radiative heating rates that closely match estimates based on Huygens measurements (Tomasko et al., 2008). Still, the shortwave heating at high zenith angles (dawn and dusk) should be regarded with caution given the plane parallel assumption, though this issue is reduced in the thick and strongly scattering lower atmosphere. Near the equator, where the Huygens and Dragonfly landing sites are located, the horizontal resolution is about 253 km in both latitude and longitude, corresponding to an area of approximately km2 per grid cell. The atmospheric vertical resolution in the lower few kilometers is maintained at several tens of meters to sufficiently resolve boundary layer variations. Subsurface layer depths follow and , where represents the diurnal skin depth for Titan and denotes the subsurface level. This configuration ensures adequate resolution for both diurnal and seasonal subsurface heat conduction by progressively increasing layer depth, with 5 layers within , 10 layers within (the annual skin depth), and 5 layers extending to the deepest region at .
The model is initialized from a cold start with uniform surface and subsurface temperatures of K, and integrated for 200 Titan years with diurnally averaged output. For diurnal analyses, the simulation is extended by one additional Titan year with output saved every 16 Earth hours (i.e., 1/24 of a Titan day), focusing on a 30 Titan-day period centered around southern midsummer in the final year.
@@ 3.1 Diurnal Variations
Figure 1 presents the diurnal evolution of surface energy balance components and surface temperature near the Huygens landing site during midsummer for the four simulation cases. In Case 4, the plains terrain type ( TIU) is assumed at the Huygens landing site (MacKenzie et al., 2019). In Case 1, the net shortwave (SW) flux increases from 0.0 to approximately 1.3 Wm-2 from nighttime to local noon, while the sensible heat flux peaks near 1.2 Wm-2 around midday, indicating that a substantial portion of incoming energy is returned to the atmosphere. This behavior reflects the limited capacity of low thermal inertia surfaces to conduct heat into the subsurface, necessitating that excess surface energy be dissipated through sensible heat flux. During the evening and nighttime, the sensible heat flux turns slightly negative, indicating rapid surface cooling and the formation of a near-surface temperature inversion. The inversion is evident in the potential-temperature structure and is discussed in Section 4.1. As thermal inertia increases (Cases 2 and 3), the daytime sensible heat flux amplitude decreases to 1.0 and 0.8 Wm-2, as subsurface heat conduction more effectively removes energy from the surface. During the evening and nighttime, the sensible heat flux shifts toward positive values, reflecting slower surface cooling. Meanwhile, the net infrared (IR) radiative flux remains relatively constant at approximately 0.07 Wm-2 across all cases, consistent with the relatively small changes in surface temperature. Case 4 shows similar behavior to Case 2 due to their comparable thermal inertias.
@@
The simulated diurnal surface temperature variations are approximately , , and K for Cases 1, 2, and 3, respectively. These values represent the mean and one standard deviation computed over a 30 Titan-day period centered on midsummer (–) in the final simulation year. For a diurnal forcing frequency of s-1 and net energy flux amplitudes of , , and Wm-2 for Cases 1, 2, and 3, respectively—derived using the same 30 Titan-day averaging method as for the simulated surface temperature amplitudes—and with an atmospheric damping coefficient of [WmK-1], obtained as the 30-day mean value, the theoretical temperature variations according to Eqs. (6) and (7) are estimated to be , 0.16, and 0.13 K. These estimates closely align with the simulated values, reinforcing the validity of the theoretical framework despite its simplification of the net energy flux as a sine-like function. For Case 4, the Wm-2 yields a theoretical expectation of K, closely matching the K as well.
@@
Despite the fact that the net energy flux—the residual of the surface energy budget—increases with thermal inertia, the amplitude of surface temperature variations decreases. The decreasing amplitude of temperature variation arises because increasing thermal inertia enhances subsurface heat conduction, effectively buffering the surface temperature despite the higher available energy. The values of corresponding to the thermal inertias used in Cases 1, 2, 3 and 4 are approximately 0.37, 1.17, 3.69 and 1.28 WmK-1, respectively. These values significantly exceed the atmospheric damping coefficient, WmK-1, indicating that subsurface heat conduction is the dominant process in regulating surface temperature variations at diurnal timescales.